# Empirically Exploring How Novices Write Software Models in Alloy

### v2

@@ 5. Threats to Validity
There exists several threats to the validity of our results. First, the dataset we use does not contain any information related to what novice users were thinking. Therefore, we are limited in some of the conclusions we can form. Second, the novice users are all master’s students from the University of Minho (UM) and the University of Porto (UP) between the fall of 2019 and the spring of 2023 semesters. While there is breadth to the date range of the study, we do not know the details of the educational performance and background of the master’s students, which may lack diversity.
We use MuAlloy to explore which submissions were almost correct. MuAlloy, and the Analyzer itself, does not support higher order quantification. Occasionally, mutating formulas produces higher order formulas that could not be solved with the current tools. These mutants are automatically pruned by MuAlloy. To limit the impact of this issues, when presenting the number of submissions that can be fixed by mutations, we include all submissions in this calculation, even if MuAlloy is unable to produce any valid mutants.
@@ 6. Related Work
Alloy4Fun. The creators of the Alloy4Fun dataset have published an experience paper outlining the first semester that they used Alloy4Fun in their classroom (alloy4fun, ). This paper focuses largely on lessons learned utilizing Alloy4Fun in the classroom, but does contain some preliminary analysis of the data collected, such as the rate of correct versus incorrect submissions and commonly reported error meesages. This data is from an early benchmark and consists of 9 models and only 5000 executions. Our our analysis is more in-depth and spans the significantly larger version of the dataset that contains 17 models and over 97,000 executions.
User Studies and Empirical Studies Over Alloy. There have been a few user studies exploring how developers work with Alloy (alloyformalise23study, ; negpos, ; nelsonuserstudy, ). The most recent study explored the debugging behavior of novice and expert users and discovered that users struggle to refine Alloy predicates using only the visual representation of scenarios (alloyformalise23study, ). Another recent study on novice Alloy users found that if users are shown a small collection of valid and invalid scenarios before writing a predicate, then users better understood what behavior the user should be trying to codify into a constraint (negpos, ). In addition, there was a user study that explored how users interact with different enumeration strategies (nelsonuserstudy, ). While our efforts focus on how to better teach and debug logical constraints, these user studies highlight best practices for how to present and guide users through the output of these constraints.
Tangent to our work, there is a static profile of 1,652 publicly available Alloy models pulled from Github that investigates how often users engage with different parts of Alloy’s grammar and explores the average size and complexity of Alloy models. This study is not concerned with the accuracy of the constraints within the model, but does give suggestions for educations based on frequently used and underutilized features. These recommendations are complementary to our recommendations.
@@ 7. Conclusion
Alloy is a commonly used modeling language that is able to represent both static and dynamic behavior of a system. Given Alloy’s popularity and well support toolset, the Analyzer, there is a growing body of work to debug faulty models and to use the Alloy in an educational setting. To help guide future research directions for both of these avenues, this paper explores a dataset of over 97,000 submissions made by novice users. We highlight several key findings, including common patterns in correct and incorrect novice submissions, how long novice users will attempt an exercise and the realistic effectiveness of existing mutant operators.

### v3

@@ 5. Threats to Validity
There exist several threats to the validity of our results. First, the dataset we use does not contain any information related to what novice users were thinking. Therefore, we are limited in some of the conclusions we can form. Second, the novice users are all master’s students from the University of Minho (UM) and the University of Porto (UP) between the fall of 2019 and the spring of 2023 semesters. While there is breadth to the date range of the study, we do not know the details of the educational performance and background of the master’s students, which may lack diversity. Third, MuAlloy, and the Analyzer itself, do not support higher-order quantification. Occasionally, mutating formulas produces higher-order formulas that can not be solved with the current tools. In theory, these higher order formula could be correct. However, since the current tools cannot evaluate them, we do not factor higher order formulas into our calculations of how many submissions are one step away.
@@ 6. Related Work
Alloy4Fun. The creators of the Alloy4Fun dataset have published an experience paper outlining the first semester that they used Alloy4Fun in their classroom (alloy4fun, ). This paper focuses largely on lessons learned utilizing Alloy4Fun in the classroom but does contain some preliminary analysis of the data collected, such as the rate of correct versus incorrect submissions and commonly reported error messages. This data is from an early benchmark and consists of 9 models and only 5000 executions. Our analysis is more in-depth and spans the significantly larger version of the dataset that contains 17 models and over 97,000 executions.
User Studies and Empirical Studies Over Alloy. There have been a few user studies exploring how developers work with Alloy (alloyformalise23study, ; negpos, ; nelsonuserstudy, ). The most recent study explored the debugging behavior of novice and expert users and discovered that users struggle to refine Alloy predicates using only the visual representation of scenarios (alloyformalise23study, ). Another recent study on novice Alloy users found that if users are shown a small collection of valid and invalid scenarios before writing a predicate, then users better understand what behavior the user should be trying to codify into a constraint (negpos, ). In addition, there was a user study that explored how users interact with different enumeration strategies (nelsonuserstudy, ). While our efforts focus on how to better teach and debug logical constraints, these user studies highlight best practices for how to present and guide users through the output of these constraints.
Tangent to our work, there is a static profile of 1,652 publicly available Alloy models pulled from GitHub that investigates how often users engage with different parts of Alloy’s grammar and explores the average size and complexity of Alloy models. This study is not concerned with the accuracy of the constraints within the model but does give suggestions for education based on frequently used and underutilized features. These recommendations are complementary to our recommendations.
@@ 7. Conclusion
Alloy is a commonly used modeling language that is able to represent static and dynamic behavior of a system. Given Alloy’s popularity and well supported toolset, there is a growing body of work to debug faulty models and to use the Alloy in an educational setting. To help guide future research directions for both of these avenues, this paper explores a dataset of over 97,000 submissions made by novice users. We highlight several key findings, including common patterns in correct and incorrect novice submissions, how long novice users will attempt an exercise, and the realistic effectiveness of existing mutant operators.